Architect Cristobal made $35,000 during the first year of his job. Each year, he received a 20% raise. What were his total earnings during the first ten years on the job?

Answer by mananth(16949)

Architect Cristobal made $35,000 during the first year of his job. Each year, he received a 20% raise. What were his total earnings during the first ten years on the job?
Sum of numbers in a G P= a1(r^n -1)/r-1
a1=35000
r=1.2
n=10
substitute with values in the formula
A= 35000*(1.2^10 -1)/(1.2-1)
=181710.774/0.2
= 908553.88

total earnings during the first ten years=$ 908553.88